### v2

# Bouncing cosmology and dynamical analysis of stability with non-minimal kinetic coupled gravity

Abstract
In this paper, we model the bounce phase, stability and the reconstruction of the universe by non-minimal kinetic coupling. In the process, we obtained importance information about the energy density and the matter pressure of the universe in relation to the previous universe through the bounce quantum phase. The novelty of the work is that the scale factor is obtained directly from the model and is fitted with an exponential function, with this view we explore the process of the early universe even the bounce phase. After that, we plot the cosmological parameters in terms of time evolution. In what follows, we investigate the stability of the model by the dynamical system analysis in a phase plane. Finally, we examine the stability of the universe, especially in the inflationary period, by using the phase space trajectories.
Non-minimal kinetic; Bouncing cosmology; Dynamical analysis of stability; The early universe.
pacs:
98.80.Cq, 04.50.Kd, 98.80.Bp
I Introduction
The importance of modified gravity theories in the study of the very universe is illustrated by the literature review of Odintsov et al. Odintsov-2023 . The review covers many aspects of modified gravity in cosmology and concentrated on the recent advances on inflationary cosmology in the modified gravity frame-work. It turns out that inflation as supported by theory and observation, is the fundamental aspects of the evolutionary characteristic of the universe, and that modified theories of gravity play a central role in this aspect because they describe able to describe a inflation field as multi fold due to its multiple short comings. Among the modified gravity theories, non-minimal gravitational theory enable one to obtain the asymmetric bounce, and do the dynamical stability analysis Banerjee et al and Odintsov et al. Banerjee-2022 ; Odintsov-2021 ; Odintsov-2020 . More relevantly to our work in this paper, Kerachian et al. Kerachian-2019 analyzed dynamics of a non-minimally coupled scalar field with an unspecified positive potential in a spatially curved FRW spacetime, and studied; its dynamics in a spatially curved FRW spacetime, defined a set of dimensionless variables which are valid for all positive, negative, and zero curvatures. The crucial point being able to define the dimensionless variables for each physical quantity so that a better interpretation of the full dynamics of the system can be attained. Further more the critical points were derived for a generic potential and singularities of the system, and then demonstrated that the determined critical points actually correspond to the de Sitter universe, the radiation universe, and the Milne universe.
Theories and observations suggest that during inflationary phase, the universe accelerated exponentially, and the result of these theories are observed as evidence for the acceleration of the universe in the late time. So that some evidence of the late universe can be related to type Ia supernovae and high redshift supernovae (Riess-1998, ; Perlmutter-1999, ; Bennett-2003, ), large scale structure Abazajian-2004 , cosmic microwave background anisotropies Spergel-2007 , Planck data Ade-2014 , and baryon acoustic oscillations Percival-2010 . During the period of inflation, the universe expanded rapidly and quickly reached a large size again. With the help of the inflation scenario, the primary problems of the universe such as the flatness and the horizon were answered. In addition to these problems, the existence of singularity is also raised as another problem of early cosmology. To analyze this problem, the universe is described as oscillatory, that is, the current universe was created due to the collapse of the previous universe. For this purpose, the Hubble parameter and the scale factor were two geometrical factors that show us the rate of expansion along the spatial direction. Now we discuss two possibilities that exist for the early time:
1) The big bang singularity occurs when the scale factor tends to zero.
2) Without reaching the singularity, the universe is increased again which represents the behavior of bouncing.
Since the scale factor can never be zero, then the space-time singularity does not occur in the early universe, as a result, the bounce scenario becomes stronger, i.e., when the scale factor reaches its minimum size, the universe begins to grow. This expression mathematically means that the slope of the scale factor or the derivative of the scale factor is zero at the mentioned point, so we call this point the bounce point. In fact, before reaching the bounce point, the previous universe contracts, and after passing this point, our new universe begins to grow again. While the size of the scale factor is decreasing at the end of the previous universe period, we can say $\dot{a}<0$, and while the size of the scale factor is increasing at the beginning of the current universe, we can say $\dot{a}>0$. From what was said, it can be concluded that the bounce occurs when the Hubble parameter is zero, and for the collapse of the previous universe and the beginning of the current universe is $H<0$ and $H>0$, respectively Cai-2014 ; Brandenberger-2012 ; Cai-2011 ; Sadeghi-2009 ; Singh-2018 ; Caruana-2020 ; Minas-2019 ; Skugoreva-2020 ; Amani-2016 ; Nojiri-2016 ; Cai-2013 ; Battefeld-2015 ; Sadeghi-2010 ; Lohakare-2022 ; Ilyas-2021 ; Ijjas-2018 ; Myrzakulov-2014 ; Shabani-2018 ; Brandenberg-2017 ; Bamba-2014 ; Tripathy-2019 ; Nojiri-2017 ; Brizuela-2010 ; Martin-Benito-2021 ; Sato-2018 ; Piao-2005 ; Cai-2009 ; Gordon-2003 ; Chattopadhyay-2017 ; Kerlick-1976 ; Alberti-2020 ; Novello-2008 ; Makarenko-2017 ; Brandenberger-2012a ; Cai-2013a ; Agrawal-2023 .
In what follows, we discuss the role of the scalar field in the early universe, which plays a key role in the description of inflationary cosmology, and it is stated that inflation is controlled by a scalar field called the inflaton field. In standard cosmology, the corresponding action is written in terms of two terms of kinetic energy and potential along with their coupling with the curvature term for the early universe. Therefore, the influence of the scalar field in the late universe and the early universe is related to the quintessence field and the inflaton field, respectively. In this case, if a non-minimal coupling between the curvature and the quintessence field is created, we will have the description of accelerated expansion Kamenshchik-2001 ; Tsujikawa-2013 , and if a non-minimal coupling between the curvature and the inflaton field is created, we will have the description of inflationary cosmology Lidsey-1991 ; Amendola-1993 ; Shokri-2021 . Also, some authors investigated another model with a newer approach called non-minimal derivative coupling Capozziello-2000 ; Sushkov-2009 ; Mostaghel-2017 . Some other authors developed the non-minimal derivative coupling model as a non-minimal kinetic coupling model, so that in this model there is a coupling between the scalar field and the kinetic energy, which means that the universe evolution scenario is described by the dominance of coupling term. The non-minimal kinetic coupling theory is a form of scalar tensor theory with a scalar field with minimal coupling to gravity. Some authors explored the theory for various aspects of cosmology in the early universe and even in the late universe Dress-1985 ; Granda-2010a ; Granda-2010b ; Granda-2010c ; Granda-2011 ; Sushkov-2012 ; Darabi-2015 ; Atazadeh-2015 ; Goodarzi-2022 ; Clifton-2012 .
Despite the fact that the non-minimal kinetic coupling model provides interesting cosmological behavior, but there is still work on the very early universe that makes us focus more on this period of the universe. Our main motivation in this study is to solve the problem of the singularity in the Big Bang by using the bouncing cosmology model. For this purpose, we want to consider the current universe as a continuation of the previous universe by the non-minimum kinetic coupling model, which is expected to bring us important information from the previous period of the universe. After that, we examine stability of the dynamical system by using the fixed points technique in phase space. Finally, by plotting phase space trajectories, we analyze the current model for stability in various modes.
This paper is organized in the following form:
In Sec. II, we present the foundation of non-minimal kinetic coupling. In Sec. III, we study the bouncing behavior and obtain the scale factor, and then calculate other cosmological parameters. In Sec. IV, we analyze the stability and instability of the dynamical system of the model and reveal the important information in the corresponding tables and graphs. Finally, in Sec. V, we provide a summary of the current job.
II Foundation of non-minimal kinetic coupling
We start the Friedmann-Robertson-Walker (FRW) metric in the following form,
$ds^{2}=-dt^{2}+a(t)^{2}\left(\frac{dr^{2}}{1-kr^{2}}+r^{2}(d\theta^{2}+\sin^{2}\theta~{}d\phi^{2})\right),$ (1)
[…]
$3H^{2}=\kappa^{2}\rho,$ (2a)
$2\dot{H}+3H^{2}=-\kappa^{2}p,$ (2b)
where $\kappa^{2}=8\pi G$, and $H=\dot{a}/a$ is the Hubble parameter. We are going to start with the following action Granda-2010a ; Granda-2010b ; Granda-2010c ; Granda-2011 , so that one has an interaction between the inflaton scalar field, $\phi$, with Ricci scalar, $R$, and Ricci tensor, $R_{\mu\nu}$,
$S=\int d^{4}x\sqrt{-g}\Big{[}\frac{1}{2\kappa^{2}}R-\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}\xi R\left(F(\phi)\partial_{\mu}\phi\partial^{\mu}\phi\right)-\frac{1}{2}\eta R_{\mu\nu}\left(F(\phi)\partial^{\mu}\phi\partial^{\nu}\phi\right)-V(\phi)\Big{]},$ (3)
where $\xi$ and $\eta$ are the coupling constants of dimensionless, $F(\phi)$ and $V(\phi)$ are an arbitrary functions of the scalar field and the system potential, respectively. We note that the inflaton field is an auxiliary scalar field that causes cosmic inflation in the early universe.
Thus, by taking variation of action (3) with respect to the metric, yields
$G_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa^{2}\left[T_{\mu\nu}^{\phi}+T_{\mu\nu}^{\xi}+T_{\mu\nu}^{\eta}\right],$ (4)
where $T_{\mu\nu}^{\phi}$, $T_{\mu\nu}^{\xi}$ and $T_{\mu\nu}^{\eta}$ are the usual energy-momentum tensor for scalar field, $\phi$, the minimally coupled terms, $\xi$ and $\eta$, respectively. The aforesaid energy-momentum tensors are written in the following form
$T_{\mu\nu}^{\phi}=\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\nabla_{\lambda}\phi\nabla^{\lambda}\phi-g_{\mu\nu}V(\phi),$ (5a)
$T_{\mu\nu}^{\xi}=\xi\Big{[}\left(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\right)\left(F(\phi)\nabla_{\lambda}\phi\nabla^{\lambda}\phi\right)+g_{\mu\nu}\nabla_{\lambda}\nabla^{\lambda}\left(F(\phi)\nabla_{\gamma}\phi\nabla^{\gamma}\phi\right)$
$-\frac{1}{2}(\nabla_{\mu}\nabla_{\nu}+\nabla_{\nu}\nabla_{\mu})\left(F(\phi)\nabla_{\lambda}\phi\nabla^{\lambda}\phi\right)+R\left(F(\phi)\nabla_{\mu}\phi\nabla_{\nu}\phi\right)\Big{]},$ (5b)
$T_{\mu\nu}^{\eta}=\eta\Big{[}F(\phi)\left(R_{\mu\lambda}\nabla^{\lambda}\phi\nabla_{\nu}\phi+R_{\nu\lambda}\nabla^{\lambda}\phi\nabla_{\mu}\phi\right)-\frac{1}{2}g_{\mu\nu}R_{\lambda\gamma}\left(F(\phi)\nabla^{\lambda}\phi\nabla^{\gamma}\phi\right)$
$-\frac{1}{2}\left(\nabla_{\lambda}\nabla_{\mu}\left(F(\phi)\nabla^{\lambda}\phi\nabla_{\nu}\phi\right)+\nabla_{\lambda}\nabla_{\nu}\left(F(\phi)\nabla^{\lambda}\phi\nabla_{\mu}\phi\right)\right)$
$+\frac{1}{2}\nabla_{\lambda}\nabla^{\lambda}\left(F(\phi)\nabla_{\mu}\phi\nabla_{\nu}\phi\right)+\frac{1}{2}g_{\mu\nu}\nabla_{\lambda}\nabla_{\gamma}\left(F(\phi)\nabla^{\lambda}\phi\nabla^{\gamma}\phi\right)\Big{]}.$ (5c)
In order to obtain the equation of motion, we take variation of action with respect to the scalar field and earn as
$-\frac{1}{\sqrt{-g}}\partial_{\mu}\Big{[}\sqrt{-g}\left(\xi RF(\phi)\partial^{\mu}\phi+\eta R^{\mu\nu}F(\phi)\partial_{\nu}\phi+\partial^{\mu}\phi\right)\Big{]}+\frac{dV}{d\phi}$
$+\frac{dF}{d\phi}\left(\xi R\partial_{\mu}\phi\partial^{\mu}\phi+\eta R_{\mu\nu}\partial^{\mu}\phi\partial^{\nu}\phi\right)=0.$ (6)
Now by inserting Eqs. (5) into Eq. (4), we obtain the modified form of the Friedmann equations, and applying the restriction on $\xi$ and $\eta$ given by $\eta+2\xi=0$, the third and forth terms of the action (3) represent a coupling the scalar field with the Einstein tensor, $G_{\mu\nu}$, that becomes $\xi G_{\mu\nu}\left(F(\phi)\partial^{\mu}\phi\partial^{\nu}\phi\right)$. In that case, by using FRW metric (1) we earn the modified Friedmann equations and the field equation in the following form
$3H^{2}=\kappa^{2}\left(\frac{1}{2}\dot{\phi}^{2}+V(\phi)+9\xi H^{2}F(\phi)\dot{\phi}^{2}\right),$ (7a)
$3H^{2}+2\dot{H}=-\kappa^{2}\left(\frac{1}{2}\dot{\phi}^{2}-V(\phi)-\xi\left(3H^{2}+2\dot{H}\right)F(\phi)\dot{\phi}^{2}-2\xi H\left(2F(\phi)\dot{\phi}\ddot{\phi}+\frac{dF}{d\phi}\dot{\phi}^{3}\right)\right),$ (7b)
$\ddot{\phi}+3H\dot{\phi}+\frac{dV}{d\phi}+3\xi H^{2}\left(2F(\phi)\ddot{\phi}+\frac{dF}{d\phi}\dot{\phi}^{2}\right)+18\xi H^{3}F(\phi)\dot{\phi}+12\xi H\dot{H}F(\phi)\dot{\phi}=0.$ (7c)
From Eqs. (2), (7a), and (7b) we write down the energy density and pressure as
$\rho=\frac{1}{2}\dot{\phi}^{2}+V(\phi)+9\xi H^{2}F(\phi)\dot{\phi}^{2},$ (8a)
$p=\frac{1}{2}\dot{\phi}^{2}-V(\phi)-\xi\left(3H^{2}+2\dot{H}\right)F(\phi)\dot{\phi}^{2}-2\xi H\left(2F(\phi)\dot{\phi}\ddot{\phi}+\frac{dF}{d\phi}\dot{\phi}^{3}\right),$ (8b)
where $\rho$ and $p$ are depend on the Hubble parameter, the scalar field, their time derivatives, and the system potential. It should be note that in the absence of coupling constant, $\xi$, we reach to standard inflaton model that when the potential energy of the inflation is larger than its kinetic energy, the negative pressure appears, so this issue can confirm the existence of energy potential in early cosmology. But in this research, in addition to potential energy being greater than kinetic energy, in the presence of coupling constant, $\xi$, the third and the forth terms of Eq. (8b) with the condition that parameters $H$, $\dot{H}$, $\phi$, $\dot{\phi}$, $\ddot{\phi}$, $F(\phi)$ and $dF/d\phi$ are positive lead to a more negative pressure, which is a convincing description for inflationary cosmology. In that case, the current model is expected to satisfy the positivity of the mentioned parameters, which will be considered in the next sections.
Therefore, the field Equation of State (EoS) reads
$\omega=\frac{p}{\rho}=\frac{\frac{1}{2}\dot{\phi}^{2}-V(\phi)-\xi\left(3H^{2}+2\dot{H}\right)F(\phi)\dot{\phi}^{2}-2\xi H\left(2F(\phi)\dot{\phi}\ddot{\phi}+\frac{dF}{d\phi}\dot{\phi}^{3}\right)}{\frac{1}{2}\dot{\phi}^{2}+V(\phi)+9\xi H^{2}F(\phi)\dot{\phi}^{2}},$ (9)
where the EoS is depend on the scalar field and its derivatives with respect to time evolution, the potential, and the Hubble parameter. In next section, we will study the behavior of bouncing universe.
III Bouncing cosmology behavior
In this section, we describe bouncing solution by non-minimal kinetic coupled reconstruction. A successful bounce requires the necessary conditions during the different phases, i.e., the contraction phase shows that the scale factor $a(t)$ is decreasing ($\dot{a}<0$), and in the expansion phase $\dot{a}>0$, but at the bouncing point, $\dot{a}=0$, and around this point $\ddot{a}>0$. Similarly, in the bouncing cosmology, the Hubble parameter, $H$, crosses zero from contraction phase ($H<0$) to expansion phase (${H}>0$), and at the bouncing point $H_{b}=0$ and $\dot{H}_{b}>0$ in which $H_{b}$ is the Hubble parameter at the bouncing point. Therefore, a successful bounce in the standard cosmology (2) has the below necessary condition around bouncing point
$\dot{H}_{b}=-\frac{\kappa^{2}}{2}(1+\omega)\rho>0,$ (10)
where the energy density is $\rho>0$, and therefore the EoS is less than $-1$ ($\omega<-1$).
In order to obtain the bouncing condition for non-minimal kinetic coupling, we sum Eqs. (8) and we obtain as
$\rho+p=\dot{\phi}^{2}+6\xi H^{2}F\dot{\phi}^{2}-2\xi\dot{H}F\dot{\phi}^{2}-2\xi H\left(2F\dot{\phi}\ddot{\phi}+\frac{dF}{d\phi}\dot{\phi}^{3}\right),$ (11)
where by applying the bouncing condition (10) we will have
$\dot{\phi}_{b}^{2}F_{b}>\frac{1}{\kappa^{2}\xi},~{}~{}~{}~{}\dot{H}_{b}F_{b}>\frac{1}{2\xi},$ (12)
where index "$b"$ is related to the bounce point. According to the bouncing theory, the big bang is the result of the beginning of an expansion period after a contraction period, so that, $t_{b}$ is the point time between contraction and expansion periods, the so-called bouncing point time.
Investigating the dynamics of a physical system such as inflationary cosmology includes terms of kinetic energy and potential energy, which the term of kinetic energy has its own common form, but the term of potential energy is very important because it determines the properties of the inflationary phase. Since the shape of the potential is very sensitive and dependent on a specific model, the chosen shape for the inflationary potential should be such that it can describe some features of inflation. Therefore, we consider the inflaton field potential as follows
$V(\phi)=\frac{\textrm{v}_{0}}{e^{\lambda\phi}+e^{-\lambda\phi}}$ (13)
where $\textrm{v}_{0}$ and $\lambda$ are constant. This type of potential ($\sim\textrm{sech}(\lambda\phi)$), which is associated with scalar fields, determines how the field evolves over time and how it affects the early expansion of the universe. In this case, the corresponding potential shape implies a scalar field that can roll down from the hilltop of near the origin and end up in the inflation period after passing through a steep region. Therefore, it could have implications for the production of perturbations in the early universe and the subsequent formation of large-scale structures. In what follows, considering the effect of the third and fourth terms of action (3) on the corresponding potential, it is necessary to consider the form of function $F(\phi)$ as a power law to make a balance between potential and non-minimal kinetic gravity. However, we consider $F(\phi)=\sum\limits_{n=0}c_{n}\phi^{2n}$ in which $c_{0}=1$, $c_{1}=c$, and $c_{n\geq 2}=0$.
Now, according to what was said above we have to solve the corresponding system. Since the solution of the present model has some complications, then we turn to the numerical solution. For this purpose, we numerically solve the system of Eqs. (7a) and (7c) for two functions of the scale factor and the scalar field in terms of cosmic time. On the other hand, the free parameters play a very important role, hence their selection is very sensitive according to the expected results. These free parameters must be chosen in such a way that the variation of the Hubble parameter in terms of cosmic time shows a transition from the negative value to the positive value with a positive slope, i.e., the bouncing solution. Therefore, we plot the graphs of cosmological parameters in terms of cosmic time with the approved values of free parameters as $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$. Fig. 1 (the solid graph) shows us that we have $\dot{a}<0$ when the universe is in the pre-big bang or contraction phase, and $\dot{a}>0$ when the universe is in the post-big bang or expansion phase. This means that the scale factor has a negative slope before the big bang and a positive slope for the expansion phase. It should be noted that the result obtained is from a numerical solution and not from an analytical solution, so it is interesting to be able to express a purely mathematical function for the scale factor in the following form
$a(t)=\frac{Ae^{2\mu t}+Be^{\mu t}+C}{\left(De^{\mu t}+1\right)^{2}},$ (14)
[…]
Figure 1: Graphs of the scale factor in terms of time evolution in the form of the numerical solution and the adopted scale factor.
Another cosmological parameters that can describe the bouncing solution is the Hubble parameter. For this purpose, we can draw the Hubble parameter from the scale factor in terms of cosmic time, with the help of the mentioned numerical solution as a solid graph shown in Fig. 2. Therefore, Fig. 2 shows us that the Hubble parameter passes from the contraction phase, $H<0$, to the expansion phase, $H>0$, with a positive slope, and we have the bouncing point ($H=0$) between two phases. In other words, from the previous universe, the value of the Hubble parameter decreases exponentially, and then when the universe approaches its bounce point it start to negatively increase exponentially till it reach zero at the bounce point (still at quantum gravity), and there after increase exponentially during the inflationary phase, and then start decrease exponentially in power tail after the inflation. Therefore, we can obtain the calculated Hubble parameter from the adopted scale factor (14) as an explicit function in terms of cosmic time in the following form
$H=\frac{2A\mu\,{e}^{2\mu t}+B\mu\,{e}^{\mu t}}{A\,{e}^{2\mu t}+B\,{e}^{\mu t}+C}-\frac{2D\mu\,{e}^{\mu t}}{D\,{e}^{\mu t}+1},$ (15)
where the corresponding coefficients are the same as the obtained values by fitting them to the adopted scale factor.
Figure 2: Graphs of the Hubble parameter in terms of time evolution in the form of the numerical solution and the calculated scale factor.
In what follows, we plot the scalar field in terms of time evolution with the same values of free parameters $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, and $\dot{\phi}(0)=0.5$, as shown in Fig. 3. Since scalar fields play an important role in cosmology, it is therefore very efficient for calculating cosmological parameters. We note that the effect of the scalar field is directly used as an auxiliary field in the calculation of cosmological parameters. Therefore, only the graph of the scalar field in terms of time evolution shown in Fig. 3 is sufficient, so that, there is no need to obtain the analytical form of its function.
Figure 3: Graph of the scalar field in terms of time evolution by $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
Next, we explore other cosmological parameters, including energy density and pressure around the bouncing point. In order to continue the numerical solution of the current system, we draw the variation of energy density and pressure resulting from Eqs. (8) in terms of time evolution as shown in Figs. 4. As it is evident from the energy density and the pressure of matter in Fig. 4, we note that the present universe entered the unstable quantum gravity phase at the bounce phase, from the stable later stage of the previous universe. Then it proceeded through an exponentially expanded phase (inflationary phase) till at about $t=5$, when it flatten out once more as in the previous universe. The pressure of the matter content of the previous universe at a late stage seems to be negatively increasing till at the bounce between $t=-5$ and $t=5$ when pressure gradient changes from negative to positive. From $t=5$ the universe seems to be decreasing quite rapidly and then it started decreasing asymptotically at about $t=10$.
Figure 4: Graphs of the energy density and the pressure in terms of time evolution by $\xi=20$, $\textrm{v}_{0}=25$, $c=1$, and $\lambda=0.2$ and with initial conditions $a(0)=1$, $\dot{a}(0)=0$, $\phi(0)=0.1$, $\dot{\phi}(0)=0.5$.
In the next section, we are going to analyse the present model in phase space by the autonomous dynamical system.
IV Dynamical system analysis
Here we explore the gravity model of non-minimal kinetic coupling by approach the dynamical system analysis. Therefore, we expect to find the stability conditions of the current system in a phase plane by using critical or fixed points. The standard form of the dynamical system is as $X^{\prime}=f(X)$, where $X$ is a column vector for appropriate auxiliary variables and function $f$ is a column vector function related to the independent equations, and the prime symbol is the derivative with respect to $N=\ln a$. Therefore, critical points $X_{c}$ satisfy condition $X^{\prime}=0$ and the linear stability properties of these critical points are determined.
For understanding of stability and instability fixed points, we put a system to an initial value that is close to its fixed point. If the trajectory of the solution of differential equation $X^{\prime}=f(X)$ comes close to this fixed point, it is called a stable fixed point, and if it moves away from this fixed point, it is called a unstable fixed point. This means that a stable fixed point can be considered as an attractor and an unstable fixed point as a repeller. A particle controlled by $X^{\prime}=f(X)$ forces the particle to move towards a stable fixed point, and an unstable fixed point will force a particle away from it.
To study the dynamics of the present system, we use the method of phase plane analysis, so we introduce the dimensionless variables as an autonomous dynamical system in the following form
$x=\frac{\kappa^{2}\dot{\phi}^{2}}{6H^{2}},~{}~{}~{}~{}~{}y=\frac{\kappa^{2}V}{3H^{2}},~{}~{}~{}~{}~{}z=6\xi H^{2}F,$ (16a)
$\chi=\frac{\kappa^{2}\dot{V}}{H^{3}},~{}~{}~{}~{}~{}\psi=6\xi H\dot{F},~{}~{}~{}~{}~{}\Lambda=\frac{\kappa^{2}\dot{\phi}\ddot{\phi}}{H^{3}},$ (16b)
where these attempt to analyze and scrutinize the model of the bouncing cosmology as an important feature. The important point that is evident here is that if the Hubble parameter becomes zero, what effect or changes would the autonomous variables have? Normally, the zero value for the Hubble parameter indicates the turning point in the early dynamics of the universe. Therefore, some autonomous variables, especially $x$ and $y$, tend to infinity and the progress of our calculations becomes undefined. To avoid this ambiguity, we should seek to limit these variables to a finite value, which in this case leads us to examine points after the neighborhood of the bounce point in the same early universe. For this reason, to advance this study, by replacing Eqs. (16) into Eqs. (7a), (7c), and (9), the following equations are obtained as
$x+y+3xz=1,$ (17a)
$\frac{\dot{H}}{H^{2}}=-\frac{\Lambda z+3\psi x+18xz+\Lambda+\chi+18x}{12xz},$ (17b)
$\omega=-\frac{2\Lambda z+6\psi x+9xz-9x+9y}{9\left(3xz+x+y\right)}-\frac{2xz}{3\left(3xz+x+y\right)}\frac{\dot{H}}{H^{2}},$ (17c)
immediately obtain $\omega$ to insert (17a) and (17b) into (17c) as follows:
$\omega=2x-\frac{\psi x}{2}-y+\frac{\Lambda y}{18x}-\frac{\Lambda}{18x}+\frac{\Lambda}{9}+\frac{\chi}{18}.$ (18)
As we know, the EoS parameter is the relationship between energy density and pressure in a dominated fluid within the universe. This parameter helps us to better understand the different periods of the universe and is introduced as an observable quantity to describe the universe. These eras include the cosmic inflation era, the radiation dominance era, the matter dominance era, the structures formation era, and dark energy dominance era. In cosmic inflation period, the universe expanded so fast that the result was a homogeneous universe. A theory of modern particle physics states that particles in high density can surprisingly overcome gravity, that is, it causes gravity to be repulsive, contrary to its attraction, and this gravitational repulsion is powerful enough, which causes rapid expansion. As a result, the EoS parameter for this period is equal to $-1$. In the period of radiation dominance, which begins after the end of the inflation period, the energy density was dominated by radiation such as photons, neutrinos, and other relativistic particles, and its EoS parameter is $1/3$. After the radiation-dominant period, the matter-dominant period appears due to the expansion and cooling of the universe, in that case, the energy density is dominated by non-relativistic matter, so that its EoS parameter is equal to zero, which is usually also referred to as cold dark matter. In the structures formation period, with the continuation of the evolution of the universe, the gravitational attraction of matter caused the formation of the structures of the universe such as galaxies, galaxy clusters, and superclusters. In the period of dark energy dominance, which is related to the late universe, it has the EoS parameter smaller than $-1/3$, so that the observational data confirms the crossing of the cosmological constant value with an EoS equal to $-1$. We note that in this work, we focus on the early periods of the universe, especially the period of inflation.
To differentiate from Eqs. (16a) with respect to $N=\ln{a}$, and to use from Eqs. (17), we earn the autonomous equations of the cosmic dynamic system as
$\frac{dx}{dN}=\frac{-3\psi x^{2}+\Lambda y-\chi x-12x^{2}+6xy-\Lambda-6x}{2(x+y-1)},$ (19a)
$\frac{dy}{dN}=-\frac{9\psi x^{2}y+2\Lambda xy-\Lambda y^{2}-2\chi x^{2}+\chi xy+36x^{2}y-18xy^{2}+\Lambda y+2\chi x+18xy}{6x(x+y-1)}.$ (19b)
In order to describe the stability conditions, we have to solve the current autonomous system by setting $f(x,y)=dx/dN=0$ and $g(x,y)=dy/dN=0$ as the fixed points ${fp}_{1}$ and ${fp}_{2}$ in which $A=\sqrt{-3\Lambda\psi-12\Lambda+6\chi+9}$ as shown in Tab. 1.
Table 1: The fixed points (Critical points).
 $Points$ ${fp}_{1}$ ${fp}_{2}$
$x$      $-\frac{\Lambda}{A+3}$    $\frac{\Lambda}{A-3}$
$y$ $-\frac{\chi}{A+3}$ $\frac{\chi}{A-3}$
Now to describe properties of the fixed points, we take linear perturbations for Eqs. (19) as $dx/dN\rightarrow dx/dN+\delta(dx/dN)$ and $dy/dN\rightarrow dy/dN+\delta(dy/dN)$. Next, we should determine the eigenvalue of these fixed points (Tab. 1). For this purpose, we write the system of differentials in the following matrix form
$J=\begin{pmatrix}\frac{\partial f}{\partial x}&\frac{\partial f}{\partial y}\\ \frac{\partial g}{\partial x}&\frac{\partial g}{\partial y}\end{pmatrix}$
where
$\frac{\partial f}{\partial x}=\frac{(3\psi+12)x^{2}+(6\psi+24)xy+(\Lambda+\chi+12)y-(6\psi+24)x-6y^{2}-\Lambda-\chi-6}{-2\left(-1+x+y\right)^{2}},$ (20a)
$\frac{\partial f}{\partial y}=\frac{x\left(3\psi x+\Lambda+\chi+18x\right)}{2\left(-1+x+y\right)^{2}},$ (20b)
$\frac{\partial g}{\partial x}=\frac{-(9\psi+54)x^{2}y^{2}+(2\Lambda+3\chi+9\psi+54)x^{2}y-\Lambda y(2xy+y^{2}-2x-2y+1)}{6x^{2}\left(-1+x+y\right)^{2}},$ (20c)
$\frac{\partial g}{\partial y}=\frac{\left(9\psi+36\right)x^{3}+\left(2\Lambda+3\chi-9\psi-18\right)x^{2}-\left(36x+2\Lambda+18y-36\right)xy-\left(\Lambda+3\chi+18\right)x-\Lambda\left(y^{2}-2y+1\right)}{-6x\left(-1+x+y\right)^{2}}.$ (20d)
Then, find the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ by setting $det(J-\lambda I)=0$ as
$\lambda_{1}=-\frac{9x^{2}\psi+2\Lambda x-\Lambda y+3\chi x+36x^{2}-18xy+\Lambda+18x}{6x\left(-1+x+y\right)},$ (21a)
$\lambda_{2}=\frac{-3x^{2}\psi-3\psi xy+6\psi x-12x^{2}-6xy+6y^{2}+\Lambda+\chi+24x-12y+6}{2\left(-1+x+y\right)^{2}},$ (21b)
where related to parameters $x$, $y$, $\chi$, $\psi$, and $\Lambda$. To substitute the amount $x$ and $y$ fixed points from Tab. 1, the eigenvalues obtain as shown in Tab. IV. We note that the our dynamical system analysis depends on the eigenvalues, which are studied as follows:
  • For real and positive eigenvalues ($\lambda_{1}>0$ and $\lambda_{2}>0$), trajectories move away from the critical point or fixed point, i.e., the track has a source, one’s introduced as an unstable node.
  • For real and negative eigenvalues ($\lambda_{1}<0$ and $\lambda_{2}<0$), trajectories move towards the critical point or fixed point, i.e., the track has a sink, one’s introduced as an stable node.
  • For real and opposite eigenvalues ($\lambda_{1}>0$ and $\lambda_{2}<0$ and vice versa), trajectories move both inwards and outward the critical point or fixed point, i.e., the trajectory simultaneously has a sink and has a source, and is introduced as the saddle point.
Table 2: Stability conditions of critical points.
{NiceTabular}
——c—c—c—c—— Fixed points Eigenvalues Acceptable conditions Stability conditions
$\lambda_{1}=\frac{A}{3}+1$ $A<-3$  (impossible)
${fp}_{1}$   Unstable
  $\lambda_{2}=\frac{A(A+3)}{\Lambda+\chi+A+3}$     $A>0,~{}\Lambda+\chi<-(A+3)$
$\lambda_{1}=-\frac{A}{3}+1$ $A>3$
${fp}_{2}$ Stable/Unstable
$\lambda_{2}=\frac{A(A-3)}{\Lambda+\chi-A+3}$ $A>3,~{}\Lambda+\chi<A-3$
$30<A<3,~{}\Lambda+\chi>A-$ 3
According to the above, we obtain acceptable conditions for eigenvalues as shown in Tab. IV. We note that $A$ is always positive, so we have $-\Lambda\psi-4\Lambda+2\chi+3>0$, it means, the obtained condition is the constraint between the parameters $\chi$, $\psi$, and $\Lambda$. However, by substituting the fixed points in Tab. 1 into Eqs. (21), the results in the second column of Tab. IV are calculated. As a result, we find acceptable conditions for fixed points $fp_{1}$ and $fp_{2}$ based on negative eigenvalues and insert the results in the third column of Tab. IV. The acceptable condition $fp_{1}$ exists only for eigenvalue $\lambda_{2}$ as $A>0$ and $\Lambda+\chi<-(A+3)$. The acceptable condition $fp_{2}$ exists for the eigenvalue $\lambda_{1}$ is only one condition in the form $A>0$, and for the eigenvalue $\lambda_{2}$ there are two conditions in the forms $A>3,\Lambda+\chi<A-3$ and $30<A<3,\Lambda+\chi>A-$ 3. We should note that in this work, we study the early universe and focus on the inflationary era immediately after the bounce.
To analyze the behavior and stability of a dynamic system, we use a phase portrait, which is a valuable tool in the study of dynamic systems. The trajectories of a dynamic system in the phase plane are represented geometrically by the phase portrait. A curve or point is drawn with the initial conditions defined for each set. For this purpose, we find the stability conditions of the inflation era according to the fourth column of Tab. IV, which we describe how to obtain it separately for different modes as follows:
The first mode: Since $A$ is always positive, so $\lambda_{1}$ is also always positive, and $\lambda_{2}$ will be negative for the acceptable condition, i.e, $A>0$ and $\Lambda+\chi<-(A+3)$. In order to represent attractor and repeller in the phase space, based on the aforesaid acceptable condition, we choose the values of $\Lambda=-3.5$, $\chi=-7$, and $\psi=1.07$, and plot phase portrait of the equations system (19) in $x-y$ plane according to the left panel of the Fig. 5. From this figure, we find that $fp_{1}$ is a saddle point or repeller, but in contrast, $fp_{2}$ is a stable point or attractor. We note that the values of EoS obtain by Eq. (18) for the fixed points as $\omega(fp_{1})\simeq-1$ and $\omega(fp_{2})\simeq 0$.
Figure 5: Left panel: Phase space trajectories on the $x-y$ plane for the values $\Lambda=-3.5$, $\chi=-7$ and $\psi=1.07$. Right panel: Phase space trajectories on the $x-y$ plane for the values $\Lambda=-2.55$, $\chi=-4.15$ and $\psi=0.55$.
The second mode: In this mode, we use the acceptable condition of $A>3$. By this condition, we choose the values of $\Lambda=-2.55$, $\chi=-4.15$, and $\psi=0.55$, and plot the equations system (19) according to the right panel of Fig. 5. The figure shows that $fp_{1}$ is a saddle point or repeller, but, $fp_{2}$ is a stable point or attractor. Also, We can obtain the values of EoS for specified fixed points as $\omega(fp_{1})\simeq-0.3$ and $\omega(fp_{2})\simeq-1$.
Figure 6: Left panel: Phase space trajectories on the $x-y$ plane for the values $\Lambda=-1.5$, $\chi=-2$ and $\psi=0.494$. Right panel: Phase space trajectories on the $x-y$ plane for the values $\Lambda=-0.2$, $\chi=-2.2$ and $\psi=3.47$.
The third mode: Herein, we consider the acceptable condition of $A>3$ and $\Lambda+\chi<A-3$ from Tab. IV. By applying this condition, we choose the values of $\Lambda=-1.5$, $\chi=-2$, and $\psi=0.494$, and plot the equations system (19) according to the left panel of Fig. 6. From this figure we find that $fp_{1}$ is a saddle point or repeller, and $fp_{2}$ is a stable point or attractor. In addition, the values of EoS obtain for the fixed points as $\omega(fp_{1})\simeq 0.1$ and $\omega(fp_{2})\simeq-1$.
The fourth mode: In this mode, considering the acceptable condition $30<A<$ 3 and $\Lambda+\chi>A-3$ in Tab. IV. This condition leads us to choose the values $\Lambda=-0.2$, $\chi=-2.2$, and $\psi=3.47$, and plot the equations system (19) according to the right panel of Fig. 6. Fig. 6 shows that both fixed points $fp_{1}$ and $fp_{2}$ are saddles or repellers. So that the EoS values for fixed points are obtained as $\omega(fp_{1})\simeq-0.68$ and $\omega(fp_{2})\simeq-1$.
In general, the aforesaid modes show an overview of the early universe, this means that after bounce point, the universe enters different periods. One of these periods is cosmic inflation with EoS equal to $-1$. For this purpose, the motivation of our choice for the values $\Lambda$, $\chi$, and $\psi$ is to establish the acceptable conditions (containing Tab. IV) and $\omega=-1$. In fact, the shape of the inflation potential has a direct effect on the stability of cosmic inflation. For this reason, we choose the potential (13) to satisfy the above description that even the phase space trajectories show the stability of the universe in the inflationary period. Therefore, the obtained results bring us a more accurate understanding of the early period of the universe.
V Conclusion
In this paper, we have modeled the evolution of the universe from the point of view of bouncing cosmology using the theory of non-minimal kinetic. For this purpose, we first addressed the issue of non-minimum kinetic theory in background flat-FRW. We restricted the minimally coupled terms $\xi$ and $\eta$ to the condition $\eta+2\xi=0$ to modify the corresponding action from the more general case to a more well-known case of the Einstein tensor. Next, we obtained the modified forms of Friedmann’s equations and the field equation, and then we acquired the energy density and the pressure of the universe.
In what follows, to investigate bouncing cosmology, we obtained the bouncing condition and we then reconstructed the evolution of the universe using the theory of non-minimal kinetic coupling. By applying the bouncing condition in the Friedman equations and the field equation, we calculated the scale factor and the Hubble parameter in terms of time evolution with numerical solution. We earned the purely mathematical function for the scale factor in the form of a exponentially function, so that the adopted scale factor was fitted with the scale factor graph points resulting from the numerical solution. In that case, we also obtained a mathematical function for the Hubble parameter from the adopted scale factor. We can see from the plot of the Hubble parameter that it passes through a contraction phase ($H<0$), to a bounce point ($H=0$) and then into an expansion phase ($H>0$), with a positive slope. Afterward, the graphs of the scalar field, the energy density, and the pressure in terms of cosmic time were plotted.
To explore the stability of the model, we used dynamical system analysis in phase plane by critical or fixed points. For the same purpose, we introduced the dimensionless variables as an autonomous dynamical system, and then we earned the autonomous equations of the cosmic dynamic system. Afterward, by setting the autonomous equations equal to zero, we obtained the fixed points according to Tab. 1, next, stability conditions of critical points were obtained according to the contents of Tab. IV. In what follows, we considered four different modes for the fixed points and then, obtained the stability points by placing constraints for each mode. After that, we drew the figures of the phase space trajectories in $x-y$ phase plane. The results obtained from these modes showed that, after the bounce point, the universe enters different periods, one of which is cosmic inflation. Since the shape of the potential has a direct effect on the stability in the inflationary period, so, choosing potential (13) brought good results for the inflationary period. Finally, the obtained results help us to better understand in the early universe. As an interesting future work, we propose the extension of the present research for a viable model.